# GME: Improving Universal Multimodal Retrieval by Multimodal LLMs

### v1

@@ C.2 Instructions
Appendix D Fused-Modal Data Synthesis Details
@@ D.1 Prompts
D.2 Filter
Two filtering methods are implemented to ensure the quality of the synthesized data. First, a text retrieval model is utilized to evaluate unreconstructed queries and their corresponding passages. We follow the framework of Promptagator (Dai et al., 2023); a query is deemed unqualified if the passage that generated it does not appear within the top 20 search results. Second, for images obtained through the Google Image Search API, we employ the CLIP model to assess image-caption relevance. Images with a relevance score below 0.2 are filtered out.
@@ D.2 Filter
| Type | Task | Query Text | Query Image | Target Text | Target Image | ||||||||||
TT |
@@
D.3 Examples of synthetic data
@@ 2. Single Language Limit
![[Uncaptioned image]](figures/dataset_examples_figures/night_q.png)
![[Uncaptioned image]](figures/dataset_examples_figures/night_c.png)
![[Uncaptioned image]](figures/dataset_examples_figures/fashion200k_c.png)
![[Uncaptioned image]](figures/dataset_examples_figures/vidore_c.png)
![[Uncaptioned image]](figures/dataset_examples_figures/visualnews_c.png)
![[Uncaptioned image]](figures/dataset_examples_figures/webqa_c.png)
![[Uncaptioned image]](figures/dataset_examples_figures/oven_q.png)
![[Uncaptioned image]](figures/dataset_examples_figures/fashioniq_q.png)
![[Uncaptioned image]](figures/dataset_examples_figures/fashioniq_c.png)
![[Uncaptioned image]](figures/dataset_examples_figures/evqa0.png)
![[Uncaptioned image]](figures/dataset_examples_figures/evqa_kb0.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/animal_wiki.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/animal_flux.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/animal_google.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/architecture_wiki.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/architecture_flux.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/architecture_google.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/artwork_wiki.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/artwork_flux.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/artwork_google.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/currency_wiki.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/currency_flux.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/currency_google.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/entertainment_wiki.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/entertainment_flux.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/entertainment_google.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/food_wiki.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/food_flux.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/food_google.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/language_wiki.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/language_flux.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/language_google.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/literature_wiki.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/literature_flux.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/literature_google.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/mythology_wiki.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/mythology_flux.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/mythology_google.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/organization_wiki.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/organization_flux.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/organization_google.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/person_wiki.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/person_flux.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/person_google.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/pharmaceutical_wiki.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/pharmaceutical_flux.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/pharmaceutical_google.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/plant_wiki.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/plant_flux.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/plant_google.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/sport_wiki.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/sport_flux.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/sport_google.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/vehicle_wiki.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/vehicle_flux.png)
@@
![[Uncaptioned image]](figures/dataset_examples_figures/vehicle_google.png)

### v2

@@ 9.2 Instructions
10 Fused-Modal Data Synthesis Details
@@ 10.1 Prompts
10.2 Filter
Two filtering methods are implemented to ensure the quality of the synthesized data. First, a text retrieval model is utilized to evaluate unreconstructed queries and their corresponding passages. We follow the framework of Promptagator [9]; a query is deemed unqualified if the passage that generated it does not appear within the top 20 search results. Second, for images obtained through the Google Image Search API, we employ the CLIP model to assess image-caption relevance. Images with a relevance score below 0.2 are filtered out.
@@ 10.2 Filter
| Type | Task | Query Text | Query Image | Target Text | Target Image | |||||||||
| Single-Modal | TT |
@@
10.3 Examples of synthetic data
@@ 2. Single Language Limit
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/night_q.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/night_c.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/fashion200k_c.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/vidore_c.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/visualnews_c.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/webqa_c.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/oven_q.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/fashioniq_q.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/fashioniq_c.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/evqa0.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/evqa_kb0.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/animal_wiki.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/animal_flux.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/animal_google.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/architecture_wiki.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/architecture_flux.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/architecture_google.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/artwork_wiki.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/artwork_flux.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/artwork_google.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/currency_wiki.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/currency_flux.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/currency_google.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/entertainment_wiki.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/entertainment_flux.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/entertainment_google.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/food_wiki.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/food_flux.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/food_google.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/language_wiki.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/language_flux.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/language_google.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/literature_wiki.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/literature_flux.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/literature_google.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/mythology_wiki.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/mythology_flux.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/mythology_google.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/organization_wiki.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/organization_flux.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/organization_google.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/person_wiki.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/person_flux.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/person_google.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/pharmaceutical_wiki.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/pharmaceutical_flux.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/pharmaceutical_google.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/plant_wiki.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/plant_flux.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/plant_google.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/sport_wiki.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/sport_flux.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/sport_google.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/vehicle_wiki.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/vehicle_flux.png)
@@
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/vehicle_google.png)